# PQPP: A Joint Benchmark for Text-to-Image Prompt and Query Performance Prediction

### v1

@@ Abstract
Text-to-image generation has recently emerged as a viable alternative to text-to-image retrieval, due to the visually impressive results of generative diffusion models. Although query performance prediction is an active research topic in information retrieval, to the best of our knowledge, there is no prior study that analyzes the difficulty of queries (prompts) in text-to-image generation, based on human judgments. To this end, we introduce the first dataset of prompts which are manually annotated in terms of image generation performance. In order to determine the difficulty of the same prompts in image retrieval, we also collect manual annotations that represent retrieval performance. We thus propose the first benchmark for joint text-to-image prompt and query performance prediction, comprising 10K queries. Our benchmark enables: (i) the comparative assessment of the difficulty of prompts/queries in image generation and image retrieval, and (ii) the evaluation of prompt/query performance predictors addressing both generation and retrieval. We present results with several pre-generation/retrieval and post-generation/retrieval performance predictors, thus providing competitive baselines for future research. Our benchmark and code is publicly available under the CC BY 4.0 license at https://github.com/Eduard6421/PQPP.
@@ 1 Introduction
Figure 1: We select a set of 10K captions from MS COCO [26] using k-means clustering. Next, we collect human relevance judgments in two scenarios: image generation and image retrieval. For each prompt/query, we generate images with two diffusion models (Stable Diffusion [36] and GLIDE [30]) and retrieve images from MS COCO with two vision-language models (CLIP [33] and BLIP-2 [24]). Based on the collected relevance judgments, we score each prompt/query in terms of generation and retrieval performance, respectively. Finally, we train and evaluate multiple prompt/query performance predictors on the proposed benchmark. Best viewed in color.
@@
In recent years, more and more people embraced the use of large language models (LLMs) instead of traditional search engines [1, 5, 53]. The advent of generative diffusion models [7, 15, 30, 36] capable of generating high-quality and realistic images triggered a similar trend in text-to-image retrieval. This paradigm shift in information retrieval (IR) calls for an extensive exploration of research topics related to generative models. For instance, query performance prediction (QPP) is an active research topic in IR  [10, 13, 18, 27, 32]. QPP refers to the process of estimating the performance level of an IR system for a given query [6]. Assessing the difficulty of a query before retrieving the results can help an IR system take informed decisions about query optimization strategies, indexing, query-specific tuning, and so on. To the best of our knowledge, there is no prior study that analyzes the difficulty of queries (prompts) in text-to-image generation and retrieval using human relevance judgments.
@@
In this study, we introduce the first dataset of prompts which are manually annotated in terms of image generation performance. We collected over 240K relevance judgments from 147 human annotators for 40K images generated by diffusion models in response to 10K prompts. We select these prompts from the Microsoft Common Objects in Context (MS COCO) dataset [26], which comprises images with captions. We chose MS COCO as a well-established data collection that allows cross-task studies and is cost-effective. It provides high-quality image annotations, as well as a wide range of everyday scenes and objects. To select a representative and diverse set of prompts from over 590K captions in MS COCO, we employ a k-means clustering algorithm on the embedding space of a sentence BERT model [42], where $k=10,000$. We extract the closest caption to each cluster center to build our set of prompts. For each prompt, we generate two images with Stable Diffusion [36] and two with GLIDE [30]. Further, we aim to study the same set of 10K prompts from two perspectives: generation performance and retrieval performance. To this end, we collect another set of 1.25M annotations from 91 human evaluators to obtain ground-truth relevance judgments for the retrieval task. We start from preliminary relevance judgments determined via the pre-trained sentence BERT [42] applied on captions. If a caption from MS COCO is similar to a query in the BERT embedding space, the image corresponding to the respective caption is added to the set of images that are potentially relevant to the respective query. Next, we collect human relevance judgments for all 10K queries. The retrieval difficulty score of each query is estimated with respect to two state-of-the-art vision-language models, CLIP [33] and BLIP-2 [24], which are tasked with retrieving images for all 10K queries. The entire process used to develop the Prompt and Query Performance Prediction (PQPP) benchmark is illustrated in Figure 1.
@@
We conduct preliminary experiments to compare prompt/query difficulty in image generation versus image retrieval. Our findings show that there is a very low correlation between the two tasks, which justifies the need to study the novel task of prompt performance prediction in image generation. We also carry out experiments with multiple pre- and post-generation/retrieval performance predictors, providing a set of competitive baselines for future research. We find that a strong supervised pre-generation/retrieval predictor is a worthy competitor for post-generation/retrieval predictors in both text-to-image generation and retrieval.
@@
  • We collect nearly 1.5M relevance judgments from human annotators to score a total of 10K queries in terms of generation and retrieval performance.
@@ 2 Related Work
Studies on QPP initially focused on textual ad hoc retrieval, where both pre- and post-retrieval predictors were considered for sparse retrieval models [8, 22, 29, 48]. Some recent studies investigated dense (neural network) retrieval models [2, 10, 11, 18, 49], as well as diverse tasks, such as conversational search [19, 27, 37]. We discuss QPP in textual ad hoc retrieval in Appendix A.4.
@@
Further contributions by Li et al. [25] delve into the challenges of estimating query difficulty with unigram language models and visual word verification, highlighting the complexities of aligning text-based queries with visual data. Meanwhile, the development of a self-supervised framework for Content-Based Image Retrieval (CBIR) systems marks a significant advancement [45], addressing the scarcity of labeled data through synthetic data and rank-based feature training. Recent efforts, such as those of Wu et al. [46] and Pavlichenko and Ustalov [31], focus on integrating human feedback into text-to-image models to refine prediction accuracy and enhance the visual appeal of generated images. Kumari et al. [23] introduced an approach to concept ablation in text-to-image synthesis, aiming to selectively prevent the generation of specific concepts.
@@
QPP evaluation. To evaluate QPP, one should determine the degree of dependence between the predicted performance and the actual performance obtained by an IR system. A correlation test serves this purpose and, for this reason, it is widely used. Some studies have highlighted the limitation of correlation coefficients for evaluating QPP in certain scenarios [17, 21, 28], and suggested alternative measures, such as the scaled Absolute Ranking Error (sARE) and its mean (sMARE) to model performance using distributions [17]. Although we agree that correlation may not be the only way to evaluate QPP, we choose to use it in the evaluation part of our study due to its wide adoption in QPP literature.
@@ 3 Proposed Benchmark
Figure 2: An example showing the annotation interface for a random prompt and the images associated with the respective prompt. For each image, the annotator can select one of the following options: high relevance, low relevance, no relevance and unrealistic. The relevance judgments of an annotator are shown for illustrative purposes. Best viewed in color.
To develop our novel Prompt and Query Performance Prediction (PQPP) benchmark, we harness the images and captions from the MS COCO [26] training set, which comprises approximately 118K images. Each of these images is accompanied by $5$ to $7$ descriptive captions, providing a rich context for our study.
To establish a unified foundation for text-to-image generation and retrieval, we select a subset of $10,000$ captions to be used across tasks either as queries for text-to-image retrieval, or as prompts for text-to-image generation. We aim for a wide variety of captions that can be clearly associated with images. The selection process entails identifying the most relevant caption for each image (from the set of captions available in MS COCO for the respective image), where the (caption, image) similarity is measured via the cosine similarity in the CLIP embedding space [33]. This reduces the number of captions to about 118K. The resulting captions are further processed by a sentence transformer based on BERT [42] to extract sentence embeddings. Aiming to enhance the distinctiveness of our final set of captions, we next apply the k-means clustering algorithm based on k-means++ initialization on the extracted embeddings. For each cluster, we select the nearest caption to the center of each cluster and include it in our final set of captions. We set $k=10,000$, thus obtaining a set of 10K prompts/queries. We next describe the procedures for collecting human relevance judgments for text-to-image generation and text-to-image retrieval in separate sections.
@@ 3.1 Prompt Performance Assessment
Generative models. To generate images for the established set of prompts, we employ the well-established Stable Diffusion [36] and GLIDE [30] generative models. For the first model, we choose the Stable Diffusion XL (base-1.0) variant, which produces images with a resolution of $1024\times 1024$ pixels. In contrast, GLIDE initially produces images at a resolution of $64\times 64$ pixels. The images are subsequently upsampled to a resolution of $256\times 256$ pixels using the upsampling model integrated in GLIDE. For each prompt, we generate two images with Stable Diffusion and two images with GLIDE.
Control and calibration. Along with the four generated images, we include the image from MS COCO associated with each prompt into the annotation process. The ground-truth image is included to calibrate user annotations (assuming that the ground-truth image should be labeled as highly relevant) or to exclude annotators that are not seriously engaged in the annotation process. To further ensure the quality of human relevance judgments, we implement a verification process based on a control set of 100 prompts. The control prompts are independently annotated by three authors of this study, following the same protocol as the other annotators. The Fleiss’ $\kappa$ coefficient among the control annotators is $0.55$. While running the full-scale annotation process, a control prompt is randomly inserted among every five prompts. This allows us to set a minimum threshold for the Cohen’s $\kappa$ coefficient to accept the relevance judgments of an annotator. The relevance judgments of an annotator are included into the study if the respective annotator (i) has at least a moderate agreement [^1] https://en.wikipedia.org/wiki/Cohen%27s_kappa#Interpreting_magnitude ($\kappa>0.4$) with each annotator in the control group, and (ii) has annotated more than $95\%$ of the ground-truth images as highly relevant.
Annotation interface. We use a custom web interface to collect relevance judgments. The prompts are presented in a random order to each annotator. Each prompt is accompanied by four generated images and one ground-truth image, which are mixed and displayed in a random order, as shown in Figure 2. For each image, annotators are given a choice among four evaluative categories: high relevance, low relevance, no relevance, and unrealistic. Participants receive thorough instructions (summarized below) before starting the annotation process. The ratings are to be assigned as follows: high relevance (score $2$) – the image depicts over half of the concepts mentioned in the prompt; low relevance (score $1$) – the image captures at least one concept, but fewer than half; no relevance (score 00) – the image is unrelated, yet realistic; unrealistic (score $-1$) – the image exhibits notable artifacts.
@@
Statistic Min Mean Max
#annotations per person 30 1,664 15,045
Fleiss’ $\kappa$ 0.41 0.54 1.00
@@
Annotation process. Our aim is to collect at least three annotations per image. In order to collect the required number of annotations (10,000 prompts $\times$ 5 images $\times$ 3 annotations = 150,000 annotations in total), we recruit $173$ annotators. The annotators providing the relevance judgments are adults having at least a bachelor or college degree. The recruited annotators willingly agreed to engage in the annotation process, after reading our terms and conditions. Annotators are allowed to opt out at any time during the annotation process. To reduce bias or uncertainty, annotators are permitted to update previously made annotations or skip specific prompts altogether. Annotators are informed about the inclusion of control prompts within their tasks, but are not given specifics on the frequency of such prompts. A fair compensation (proportional to the number of annotated prompts) is given to each annotator with a Cohen’s $\kappa$ coefficient higher than $0.4$ on the control prompts. Based on our selection criteria, $26$ annotators are excluded from the process. This leaves us with $147$ valid annotators. Given the asynchronous nature of the annotation process, several prompts ended up having more than three annotations, leading to a total of 244,650 annotations. When a prompt has more than three relevance judgments, we keep the annotations provided by the top three annotators with the highest Cohen’s $\kappa$ coefficients (with respect to the control prompts). A brief set of statistics about the enrolled annotators is presented in Table 1. Notably, the Fleiss’ $\kappa$ coefficient computed across all annotators is consistent with that of the control annotators. Upon excluding the annotations corresponding to the ground-truth images, we find that most images are voted as highly relevant (see Figure 4 in Appendix A.5), confirming that Stable Diffusion and GLIDE generally produce relevant results.
Measuring prompt performance. To derive the final prompt performance in image generation, we first map the relevance categories to numerical values, as follows: high relevance is mapped to 2, low relevance to 1, no relevance to 0, and unrealistic to -1.
To exclude outlier annotations, we group the four relevance categories into two high-level categories. The first category, combining the high relevance and low relevance annotations, represents images that are at least somewhat relevant to the prompt. The second category, combining the unrealistic and no relevance labels, represents images that are not acceptable for the given prompt, either because they are irrelevant or unrealistic. We hereby acknowledge that the distinction between high relevance and low relevance is more difficult to determine, involving a fine assessment of how many of the prompt elements are depicted in the image. This requires evaluators to consider not just the presence of these elements, but also their significance and portrayal within the image, making the distinction between high and low relevance inherently more subjective and challenging. In contrast, the distinction between the high-level categories (relevant and irrelevant) can be easily assessed.
@@
We employ a majority voting mechanism on the high-level categories to decide if a generated image is either relevant or irrelevant. Since there are only two high level categories and three annotations per image, there is no need to break ties (a majority always exists). The majority voting is performed to rule out outlier annotations. The final relevance of an image is given by averaging the scores (between $-1$ and $2$) associated with the votes forming the majority. The performance of a prompt is given by the average relevance score computed across the four generated images. We further refer to the resulting score as human-based prompt performance (HBPP).

### v2

@@ Abstract
Text-to-image generation has recently emerged as a viable alternative to text-to-image retrieval, driven by the visually impressive results of generative diffusion models. Although query performance prediction is an active research topic in information retrieval, to the best of our knowledge, there is no prior study that analyzes the difficulty of queries (referred to as prompts) in text-to-image generation, based on human judgments. To this end, we introduce the first dataset of prompts which are manually annotated in terms of image generation performance. Additionally, we extend these evaluations to text-to-image retrieval by collecting manual annotations that represent retrieval performance. We thus establish the first joint benchmark for prompt and query performance prediction (PQPP) across both tasks, comprising over 10K queries. Our benchmark enables (i) the comparative assessment of prompt/query difficulty in both image generation and image retrieval, and (ii) the evaluation of prompt/query performance predictors addressing both generation and retrieval. We evaluate several pre- and post-generation/retrieval performance predictors, thus providing competitive baselines for future research. Our benchmark and code are publicly available at https://github.com/Eduard6421/PQPP.
@@ 1 Introduction
Figure 1: We select a set of 10K captions from MS COCO [30] using k-means clustering, which are further merged with prompts from DrawBench [41]. Next, we collect human relevance judgments in two scenarios: image generation and image retrieval. For each prompt/query, we generate images with two diffusion models (Stable Diffusion XL [35] and GLIDE [33]) and retrieve images from MS COCO with two vision-language models (CLIP [37] and BLIP-2 [27]). Based on the collected relevance judgments, we score each prompt/query in terms of generation and retrieval performance, respectively. Finally, we train and evaluate multiple prompt/query performance predictors on the proposed benchmark. Best viewed in color.
@@
In recent years, more and more people embraced the use of large language models (LLMs) instead of traditional search engines [1, 7, 62]. The advent of generative diffusion models [9, 18, 33, 40] capable of generating high-quality and realistic images triggered a similar trend in text-to-image retrieval. This paradigm shift in information retrieval (IR) calls for an extensive exploration of research topics related to generative models. One is query performance prediction (QPP), which refers to the process of estimating the performance level of an IR system for a given query [8], this being an active research topic in IR [13, 16, 20, 31, 36]. In the context of generative AI for text-to-image generation, if a prompt is predicted to be difficult, the system could initiate a conversation to refine the prompt in order to overcome the difficulty and improve the final output. Moreover, the system could indicate to the user its inability to provide a satisfactory image, or it could give positive feedback to the user when the prompt is predicted as easy. In general, when a prompt/query is predicted as difficult for a generation/retrieval system, additional processes can be activated, such as the automatic reformulation, the allocation of extra resources, or the addition of pre- or post-processing. We discuss further applications and concrete use cases of prompt performance prediction in Appendix 8.
With the growing popularity of generative models and retrieval augmented generation, understanding prompt effectiveness in both generation and retrieval becomes increasingly essential. While text-to-image generation and retrieval share common elements, they differ fundamentally in terms of task requirements and success criteria. In text-to-image generation, prompts must guide models to produce images that not only capture the specified elements, but also meet aesthetic and contextual expectations. Conversely, in retrieval, the query role is to retrieve pre-existing images that match the query as closely as possible within a given dataset, where success depends on how well the query aligns with the available data. Studying both tasks together is critical. Examining prompt/query difficulty across generation and retrieval allows us to determine whether the factors that make a prompt challenging in generation (e.g. complexity, specificity, ambiguity) are the same as those in retrieval. To the best of our knowledge, there is no prior study that jointly analyzes the difficulty of prompts/queries in text-to-image generation and retrieval using human relevance judgments. In this study, we introduce the first dataset of prompts/queries which are manually annotated in terms of both image generation performance and image retrieval performance. Additionally, we provide baseline results with various pre- and post-performance predictors, setting a standard for future multimodal QPP research.
Our new Prompt and Query Performance Prediction (PQPP) benchmark comprises 10,200 text samples that serve as prompts for text-to-image generation and queries from text-to-image retrieval. To obtain representative prompts/queries for both generation and retrieval, we combine the 200 prompts from DrawBench [41] with 10,000 diverse captions from the Microsoft Common Objects in Context (MS COCO) dataset [30]. While DrawBench is a recently proposed set of prompts that aims to test the capability of generative models in diverse scenarios, MS COCO is a well-established data collection that allows cross-task studies and is cost-effective. MS COCO provides high-quality image annotations, as well as a wide range of everyday scenes and objects, being widely used across several computer vision tasks, e.g. object detection, object segmentation, and image captioning. Notably, with the advent of text-to-image generation, MS COCO has also been employed in many works for the text-to-image generation task [5, 24, 41, 55]. To select a representative and diverse set of prompts from over 590K captions in MS COCO, we employ a k-means clustering algorithm on the embedding space of a sentence BERT model [47], where $k\!=\!10,\!000$.
@@
To analyze prompt performance in both generation and retrieval, we examine the same set of 10,200 prompts across two generative models and two retrieval models (see Figure 1). For the generation task, we use SDXL [35] and GLIDE [33], each generating two images per prompt. Both models are state-of-the-art diffusion models, yet they rely on different designs, which leads to variability in terms of image quality and relevance. Next, we collect over 247K relevance judgments from 147 human annotators for 40,800 images generated by the chosen diffusion models. For the retrieval task, we employ two state-of-the-art vision-language models, CLIP [37] and BLIP-2 [27], to retrieve images for the 10,200 queries. These models use different architectures and training data, impacting their retrieval performance for the same query. We start from preliminary relevance judgments determined via the pre-trained sentence BERT [47] applied on captions. If a caption from MS COCO is similar to a query in the BERT embedding space, the image corresponding to the respective caption is added to the set of images that are potentially relevant to the respective query. Next, we collect 1.39M ground-truth relevance judgments from 91 human evaluators to annotate the 10,200 queries. In summary, we collect over 1.6M annotations to estimate the difficulty score of each query included in PQPP, from both generation and retrieval perspectives.
@@
We conduct preliminary experiments to compare prompt/query difficulty in image generation vs. image retrieval. Our findings show that there is a very low correlation between the two tasks (see Section 4), which justifies the need to study the novel task of prompt performance prediction in image generation. We also carry out experiments with multiple pre- and post-generation/retrieval performance predictors, providing a set of competitive baselines for future research. We find that a strong supervised pre-generation/retrieval predictor is a worthy competitor for post-generation/retrieval predictors in both text-to-image generation and retrieval. To further demonstrate the usefulness of PQPP, we carry out cross-model, cross-dataset and cross-task experiments, evaluating the generalization capabilities of performance predictors in challenging settings.
@@
  • We collect over 1.6M relevance judgments from human annotators to score a total of 10,200 queries in terms of generation and retrieval performance.
@@ 2 Related Work
Studies on QPP initially focused on textual ad hoc retrieval, where both pre- and post-retrieval predictors were considered for sparse retrieval models [11, 23, 32, 56]. Some recent studies investigated dense (neural network) retrieval models [3, 13, 14, 20, 57], as well as diverse tasks, such as conversational search [21, 31, 42]. We discuss QPP in textual ad hoc retrieval in Appendix 9.
@@
Further contributions by Li et al. [28] delve into the challenges of estimating query difficulty with unigram language models and visual word verification, highlighting the complexities of aligning text-based queries with visual data. Meanwhile, the development of a self-supervised framework for Content-Based Image Retrieval (CBIR) systems, addressing the scarcity of labeled data through synthetic data and rank-based feature training, marks a significant advancement [50]. Recent efforts, such as those of Wu et al. [52] and Pavlichenko and Ustalov [34], focus on integrating human feedback into text-to-image models to refine prediction accuracy and enhance the visual appeal of generated images. Kumari et al. [26] introduced an approach to concept ablation in text-to-image synthesis, aiming to selectively prevent the generation of specific concepts.
@@
Human feedback in text-to-image generation. Human feedback can provide fundamental insights for text-to-image generators, which can harness the provided annotations to increase the quality of the generated images. To this end, several works [25, 29, 52, 54] collected human feedback for generated images and developed novel benchmarks to train and evaluate generative models. Kirstain et al. [25] obtained preference annotations by asking humans to choose an image from a pair of generated images. Wu et al. [52] also collected a dataset of human preferences on generated images. The dataset was used to train a classifier to output a human preference score. The classifier was further used to fine-tune Stable Diffusion. Xu et al. [54] asked workers to rank several generated images and rate them based on their quality. The annotations were used to train a reward model, which provided feedback for training diffusion models. Unlike previous studies, Liang et al. [29] collected rich annotations consisting of fine-grained scores, implausible regions and misaligned words. The annotations were further used to improve image generation.
While we recognize the outstanding merits of previous studies that collected human feedback for improving text-to-image generation [25, 29, 52, 54], we would like to highlight that the focus of our work is different. To our knowledge, our work represents the first attempt to study the task of predicting the performance of text-to-image generators on a given prompt.
@@ 3 Proposed Benchmark
Figure 2: An example showing the annotation interface for a random prompt and the images associated with the respective prompt. For each image, the annotator can select one of the following options: high relevance, low relevance, no relevance and unrealistic. The relevance judgments of an annotator are shown for illustrative purposes. The locations of images are randomly generated, each time they are displayed. Hence, the annotators do not see the images in the same order, which eliminates positional biases. Best viewed in color.
To develop our novel Prompt and Query Performance Prediction (PQPP) benchmark, we first harness the images and captions from the MS COCO [30] training set, which comprises approximately 118K images. Each of these images is accompanied by $5$ to $7$ descriptive captions, providing a rich context for our study.
To establish a unified foundation for text-to-image generation and retrieval, we select a subset of $10,000$ captions to be used across tasks either as queries for text-to-image retrieval, or as prompts for text-to-image generation. We aim for a wide variety of captions that can be clearly associated with images. The selection process entails identifying the most relevant caption for each image (from the small set of captions available in MS COCO for the respective image), where the (caption, image) similarity is measured via the cosine similarity in the CLIP embedding space [37]. This reduces the number of captions to about 118K. The resulting captions are further processed by a sentence transformer based on BERT [47] to extract sentence embeddings. Aiming to enhance the distinctiveness of our final set of captions, we next apply the k-means clustering algorithm based on k-means++ initialization on the extracted embeddings. For each cluster, we select the nearest caption to the center of each cluster and include it in our final set of captions. We set $k\!=\!10,\!000$ to obtain 10K prompts/queries. Although MS COCO captions have been used before to train text-to-image generators [5, 24, 41, 55], prompts written by users can often be more descriptive [10]. To this end, we add all the prompts from DrawBench [41] to our set, resulting in a combined set of 10,200 prompts/queries. We next describe the procedures for collecting human relevance judgments on both tasks.
@@ 3.1 Prompt Performance Assessment
Generative models. To generate images for our set of prompts, we employ the well-established SDXL [35] and GLIDE [33] generative models. For the first model, we choose the Stable Diffusion XL (base-1.0) variant, which produces images with a resolution of $1024\times 1024$ pixels. In contrast, GLIDE initially produces images at a resolution of $64\times 64$ pixels. The images are subsequently upsampled to a resolution of $256\times 256$ pixels using the upsampling model integrated in GLIDE. For each prompt, we generate two images with SDXL and two images with GLIDE.
Control and calibration. Along with the four generated images, we include the image from MS COCO associated with each prompt into the annotation process. The ground-truth image is included to calibrate user annotations (assuming that the ground-truth image should be labeled as highly relevant) or to exclude annotators that are not seriously engaged in the annotation process. To further ensure the quality of human relevance judgments, we implement a verification process based on a control set of 100 prompts. The control prompts are independently annotated by three authors of this study, following the same protocol as the other annotators. The Fleiss’ $\kappa$ coefficient among the control annotators is $0.55$. While running the full-scale annotation process, a control prompt is randomly inserted among every five prompts. This allows us to set a minimum threshold for the Cohen’s $\kappa$ coefficient to accept the relevance judgments of an annotator. The relevance judgments of an annotator are included into the study if the respective annotator (i) has at least a moderate agreement [^1] Wiki article on Cohen’s $\kappa$: Interpreting Magnitude ($\kappa>0.4$) with each annotator in the control group, and (ii) has annotated more than $95\%$ of the ground-truth images as highly relevant.
Annotation interface. We use a custom web interface to collect relevance judgments. The prompts are presented in a random order to each annotator. Each prompt is accompanied by four generated images and one ground-truth image, which are mixed and displayed in a random order, as shown in Figure 2. For each image, annotators are given a choice among four evaluative categories: high relevance, low relevance, no relevance, and unrealistic. Participants receive thorough instructions (summarized below) before starting the annotation process. The ratings are to be assigned as follows: high relevance (score $2$) – the image depicts over half of the concepts mentioned in the prompt; low relevance (score $1$) – the image captures at least one concept, but fewer than half; no relevance (score 00) – the image is unrelated, yet realistic; unrealistic (score $-1$) – the image exhibits notable generation artifacts.
Statistic Min Mean Max
#annotations per person 30 1,681 15,845
Fleiss’ $\kappa$ 0.41 0.54 1.00
@@
Annotation process. Our aim is to collect at least three annotations per image. In order to collect the required number of annotations (10,200 prompts $\times$ 5 images $\times$ 3 annotations = 153,000 annotations in total), we recruit $173$ annotators. Based on our selection criteria, $26$ annotators are excluded from the process. This leaves us with $147$ valid annotators. Given the asynchronous nature of the annotation process, several prompts ended up having more than three annotations, leading to a total of 247,050 annotations. When a prompt has more than three relevance judgments, we keep the annotations provided by the top three annotators with the highest Cohen’s $\kappa$ coefficients (with respect to the control prompts). Some statistics about the enrolled annotators are presented in Table 1. Notably, the Fleiss’ $\kappa$ coefficient computed across all annotators is consistent with that of the control annotators. Upon excluding the annotations corresponding to the ground-truth images, we find that most images are voted as highly relevant (see Figure 7 in Appendix 10), confirming that SDXL and GLIDE generally produce relevant results.
@@
Measuring prompt performance. To derive the final prompt performance in image generation, we first map the relevance categories to numerical values, as follows: high relevance is mapped to 2, low relevance to 1, no relevance to 0, and unrealistic to -1. To exclude outlier annotations, we group the four relevance categories into two high-level categories. The first category, combining the high relevance and low relevance annotations, represents images that are at least somewhat relevant to the prompt. The second category, combining the unrealistic and no relevance labels, represents images that are not acceptable for the given prompt, either because they are irrelevant or unrealistic. We employ a majority voting mechanism on the high-level categories to decide if a generated image is either relevant or irrelevant. The Fleiss’ $\kappa$ coefficient for these categories is $0.75$, suggesting that the annotations are more consistent at this coarse level. Since there are only two high level categories and three annotations per image, there is no need to break ties (a majority always exists). The majority voting is performed to rule out outlier annotations. The final relevance of an image is given by averaging the scores (between $-1$ and $2$) associated with the votes forming the majority. The performance of a prompt is given by the average relevance score computed across the generated images. We further refer to the resulting score as human-based prompt performance (HBPP).
More details about the prompt performance assessment are discussed in Appendix 10.